OpenStudy (mathlegend):

Write the equation of the line in slope-intercept form. Parallel to 2x-3y= 6 and passing through (-2,5)

OpenStudy (mathlegend):

What I had did here was turned the problem into slope intercept form

OpenStudy (anonymous):

then use the slope-intercept form \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

sorry, it's point-slope form

OpenStudy (mathlegend):

So do I use 2 for the slope or \[\frac{ 2}{ 3}\]

OpenStudy (mathlegend):

I used 2

OpenStudy (mathlegend):

I think I was supposed to use the other number .... oops...

OpenStudy (mathlegend):